OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 4x - y - 2z if x = -2, y = 3, and z = -2 -15 15 -7 7

OpenStudy (anonymous):

4(-2) - 3 - 2(-2) = -8 -3 - (-4) = -7